ADNAN MUHAMMED

Grade 12,

if 4 sinB=3 sin(2A+B) prove that 7 cot(A+B)=cot A.solve it fast

if 4 sinB=3 sin(2A+B) prove that 7 cot(A+B)=cot A.solve it fast
 

Grade:10

1 Answers

Arun
5 years ago
sin (2A + B) / sinB = 4/3
 
sin ((A+B )+ A) / sin ((A + B) – B) = 4/3
 
now 
 
[sin (A + B) cos A + Cos (A + B) sin A ] / [sin (A+ B) cosB – cos(A+B) sinB] = 4/3
 
apply componendo ad dividendo rule and you will surely get your answer